HOPING TO GET SOME HELP WITH THIS QUESTION!

Ryan Corporation has the following information relating to its manufacturing and selling of cakes:

Average current selling price: $14.00

Average variable cost of cakes: $6.00

Fixed yearly costs: $240,000

The income tax rate is 35%

If Ryan Corporation wanted to make a net income after tax of $100,000, how many cakes would they have to produce and sell?

a. 49,231 cakes

b. 45,333 cakes

c. 54,443 cakes

d. 57,300 cakes
Answer #1

Let the number of units be A

then total revenue = 14*A

total cost = 6*A+240000

Income before tax = TR-TC

As 65%(100-35%) of income is actual profit after tax which should be equal to 100000

So, (14*A- (6*A+240000)) x .65 = 1000000

A= (8A-240000) x .65 = 100000

A= 256000/8*.65

A=49230.76

A=49231 cakes

Option(A)